Manufacturers of electrical equipment produce incandescent lamps with a lifespan of 5000 hours. If the standard deviation is 60 hours. What is the probability that an incandescent lamp will last between 5000 and 5100 hours?


1
Expert's answer
2021-12-29T11:41:37-0500

"P(5000\leq x\leq 5100)\\=P(\frac{5000-5000}{60} \leq z \leq \frac{5100-5000}{60})\\
P(0\leq z\leq 1.67)\approx 0.4525"